OpenStudy (anonymous):

Find all real number solutions to the system of equations: y=e^(4x) y=e^(2x)+6

OpenStudy (nikvist):

\[e^{4x}=e^{2x}+6\]\[e^{4x}-e^{2x}-6=0\]\[(e^{2x}+2)(e^{2x}-3)=0\]\[e^{2x}=3\quad\Rightarrow\quad y=9\]\[x=\ln{\sqrt{3}}\]

OpenStudy (anonymous):

e^4x = e ^ 2x + 6 e^4x - 6 ^2x - 6 = 0 ( e^2x + 2) (e^2x - 3) = 0 e^2x = 3 > y = 9 x = In √3